A 96.2 cm long straight section of wire is located entirely inside a uniform magnetic field of |B| = 0.650 T. The wire is perpendicular to the direction of the magnetic field. When a current runs through the wire a magnetic force of |F| = 0.630 N acts on that section of wire. Calculate the size of the current.
Answer #1

Given:

Length of wire, L = 96.2 cm = 0.962 m

Magnetic field, |B| = B = 0.650 T

Magnitude of Force = |F| = F = 0.630 N

To find : current, i = ?

Using the formula

F = i*L*B

Putting the values in above formula,

0.630 N = i*(0.962 m)*(0.650 T)

Or, 0.630 N = i*(0.6253) T m

So the current in the wire, i = 1.0075 A

In case of any doubt please discuss in comment section below.